OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)? 4 - -4

OpenStudy (anonymous):

ur equation is y2-y1/x2-x1 the resulting anser is a fraction and is the slope

OpenStudy (anonymous):

is it -1/4 or 1/4

OpenStudy (anonymous):

slope =change in y/change in x dY/dX=y2-y1/x2-x1

OpenStudy (anonymous):

is it negative or positive?

OpenStudy (anonymous):

->(-3-(-5)/(7-(-1) =-3+5/7+1

OpenStudy (anonymous):

What grade are you?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

I'm hoping you can do the rest

OpenStudy (anonymous):

Telling you what the answer is doesn't help you learn anything